The distance between Hazleton and Pleasantville is 100miles. At 2:00 p.m. Mike starts riding his bicycle from Hazleton to Pleasantville at a constant speed of 18 mi/h. Twenty minutes later Katie leaves Hazleton for Pleasantville traveling at a rate of 30 mi/h. How far from Pleasantville are they when Katie catches up with Mike?
Answer by mananth(16946)

Let the distance from Hazelton when they meet be x
Let time taken by Mike be y
speed mike =18m/h
Distance = speed * time
x= 18y
Katie speed = 30m/h
time = y-(1/3) hours
distance =x
x= (y-1/3) *30
=((3y-1)/3 )*30
equate both
18y=30y-10
-12y= -10
y= 5/6 hours
So Mike takes 5/6 hours
x=18y
x= 18*5/6
=15 miles
so they are 100-15 = 85 miles from Pleasantville